x-6y=31
7y-8x=2
whatn is oordered pair

Answer by stanbon(75887)

x-6y=31
7y-8x=2
-------
Rearrange:
x = 6y+31
----
Substitute:
7y - 8(6y+31) = 2
------
7y -48y - 248 - 2 = 0
-41y = 250
y = -250/41
-----
Solve for "x":
x = 6y+31
x = 6(-250/41) + 31
----
x = -1500/41 + 31
x = -229/41
-------------
Cheers,
stan H.
---------------